solve and check each equation or inequality.
w-3/4=7/12

Answer by JulietG(1812)

First, make each fraction use the same denominator. Since 4 goes into 12, we'll use that.
The problem now becomes:
w - 9/12 = 7/12 (9/12 is the same as 3/4)
Add 9/12 to each side
w = 16/12
Simplify
w = 1 4/12
Simplify further
w = 1 1/3